### v3

# From Sorting Algorithms to Scalable Kernels: Bayesian Optimization in High-Dimensional Permutation Spaces

Abstract
Bayesian Optimization (BO) is a powerful tool for black-box optimization, but its application to high-dimensional permutation spaces is severely limited by the challenge of defining scalable representations. The current state-of-the-art BO approach for permutation spaces relies on an exhaustive $\Omega(n^{2})$ pairwise comparison, inducing a dense representation that is impractical for large-scale permutations. To break this barrier, we introduce a novel framework for generating efficient permutation representations via kernel functions derived from sorting algorithms. Within this framework, the Mallows kernel can be viewed as a special instance derived from enumeration sort. Further, we introduce the Merge Kernel , which leverages the divide-and-conquer structure of merge sort to produce a compact, $\Theta(n\log n)$ to achieve the lowest possible complexity with no information loss and effectively capture permutation structure. Our central thesis is that the Merge Kernel performs competitively with the Mallows kernel in low-dimensional settings, but significantly outperforms it in both optimization performance and computational efficiency as the dimension $n$ grows. Extensive evaluations on various permutation optimization benchmarks confirm our hypothesis, demonstrating that the Merge Kernel provides a scalable and more effective solution for Bayesian optimization in high-dimensional permutation spaces, thereby unlocking the potential for tackling previously intractable problems such as large-scale feature ordering and combinatorial neural architecture search.
1 Introduction
As one of the most widely adopted approaches to black-box optimization, Bayesian optimization (BO) Shahriari et al. (2015) has found broad application in machine-learning hyper-parameter tuning Wu et al. (2019), financial portfolio optimization Gonzalvez et al. (2019), chemical and material discovery Luo et al. (2025), and catalyst formulation design Xie et al. (2023). BO employs probabilistic surrogate models—most commonly Gaussian processes (GPs)—to approximate the unknown objective and uses an acquisition function to balance exploration and exploitation, thereby approaching the global optimum with the fewest possible evaluations.
Most research on BO to date has concentrated on continuous Greenhill et al. (2020) and categorical design spaces Garrido-Merchán & Hernández-Lobato (2020); Nguyen et al. (2020); Bartoli et al. (2025), whereas applications to permutation spaces remain comparatively underexplored. This gap is striking given the ubiquity of permutation optimization in both theory and practice: canonical examples include the traveling-salesperson problem (TSP), the sequencing of operations in automated experimental pipelines Guidi et al. (2020), and the sequential order-of-addition experiments Lin & Rios (2025). In addition, permutation optimization frequently arises in diverse AI applications, including scheduling tasks in robotic planning Alatartsev et al. (2015), optimizing experimental sequences Blau et al. (2022), and other sequential decision-making problems Sun & Giles (2001); Wen et al. (2023). Therefore, extending Bayesian optimization to permutation spaces holds significant theoretical and practical value for artificial intelligence as well as broader academic and industrial applications.
[…]
To address the limitations of existing approaches, we propose a sorting-algorithm–driven kernel-design framework for permutations and instantiate it with the Merge kernel, which reduces the feature dimension to $O(n\log n)$—the information-theoretic lower bound for encoding a permutation. The central insight is that any comparison-based sorting algorithm is defined by a fixed sequence of element comparisons; recording the binary outcome of each comparison yields a feature vector for the permutation. Choosing an algorithm with a deterministic comparison tree—such as merge sort or bitonic sort—thus produces a representation that is both fixed in length and highly compact.
Contributions. Our work makes three principal contributions:
  • General framework. We propose a unified design framework that constructs permutation-space kernels by treating any comparison-based sorting algorithm as a feature generator. Within this view, the classic Mallows kernel is recovered as the special case obtained when the framework is instantiated with enumeration sort.
  • Merge kernel. Applying the framework to merge sort yields Merge Kernel, whose $O(n\log n)$ construction matches the information-theoretic lower bound on comparison complexity.
  • Comprehensive evaluation. We assess the effectiveness of our kernels on diverse synthetic and real-world benchmarks. Results on low-dimensional benchmarks show competitive performance against the state-of-the-art Mallows kernel, while it significantly outperforms the Mallows kernel on high-dimensional benchmarks.
Our results demonstrate that the Merge kernel provides a practical and efficient tool for permutation optimization, significantly enhancing BO’s applicability to diverse AI scenarios.
2 Background and Related Works
2.1 Permutation Optimization
Here we describe the problem formulation of permutation optimization with a fixed length $n\in\mathbb{N}$. Let $n]=\{1,2,...,n\}$, a permutation is a function $\pi:[n]\longrightarrow[n]$ such that $\pi$ is bijective. The set of all permutations of $[n]$ is the symmetric group
$\mathcal{S}_{n}\;=\;\bigl\{\pi\mid\pi:[n]\longrightarrow[n]\text{ is bijective}\bigr\}.$
We are given a costly-to-evaluate, possibly noisy black-box function $f:\;\mathcal{S}_{n}\longrightarrow\mathbb{R}$, which assigns a real-valued quality (e.g., cost, loss, reward) to every permutation $\pi$. Hence, the optimization problem can be formulated as
$\pi^{*}\;=\;\text{arg min}_{\pi\;\in\;\mathcal{F}}f(\pi)$
where $\mathcal{F}\subseteq\mathcal{S}_{n}$ is the feasible set. In this study we only consider the unconstrained case, therefore we have $\mathcal{F}\;=\;\mathcal{S}_{n}$; in practice $\mathcal{F}$ may exclude permutations violating domain rules.
2.2 Bayesian Optimization
Bayesian optimization (BO) Shahriari et al. (2015) is an optimization algorithm for black-box objective functions that no closed-form expression or gradient information is available and whose evaluation is often an expensive physical or computational experiment. The algorithm first fits the observed data with a surrogate model, most commonly a Gaussian process (GP) Williams & Rasmussen (2006), and then employs an acquisition function to select the next query point, balancing exploration of uncertain regions against exploitation of promising areas identified by the surrogate.
The kernel $K(\mathbf{x},\mathbf{x}^{\prime})$ is the central design lever in a Gaussian-process surrogate: it defines the similarity metric between inputs, thereby specifying the prior smoothness assumptions and, through GP inference, the posterior mean and uncertainty. Extending BO to any new search domain is therefore tantamount to endowing that domain with an appropriate kernel function. Whereas Euclidean spaces typically rely on Gaussian (RBF) kernels, discrete structures—and permutations in particular—require bespoke constructions that faithfully encode ordering relationships. For example, Oh et al. (2022) uses the Position kernel Zaefferer et al. (2014) and L-ensemble with Acquisition Weights to extend BO on permutation space to a batched BO scheme.
While significant progress has been made in scaling BO to high-dimensional continuous domains (e.g., Eriksson et al. (2019)Wang et al. (2016)), extending BO to large-scale structured discrete domains like permutations presents a distinct set of challenges centered on kernel design, which is the primary focus of this work.
2.3 Mallows Kernel for Permutation Space
BOPS-H Deshwal et al. (2022) is the current state-of-the-art BO algorithm for permutation optimization, which proposes to employ Mallows kernel Jiao & Vert (2015) on the symmetric group $\mathcal{S}_{n}$ in a similar manner to the RBF kernel on the Euclidean space. The Mallows kernel $K_{mal}(\pi,\pi^{\prime})$ for the permutation pair ($\pi$, $\pi^{\prime}$) is defined as the exponential negative of the number of discordant pairs $n_{d}(\pi,\pi^{\prime})$ between $\pi$ and $\pi^{\prime}$:
$K_{Mal}(\pi,\pi^{\prime})=\exp(-ld(\pi,\pi^{\prime}))$ (1)
where $l\geq 0$ is the length-scale parameter of the Mallows kernel, and $d(\pi,\pi^{\prime})$ is the Kendall-$\tau$ distance Kendall (1938) which counts the number of pairs of elements ordered oppositely by $\pi$ and $\pi^{\prime}$:
$d\left(\pi,\pi^{\prime}\right)=\sum_{i<j}$ $\left[1_{\pi(i)>\pi(j)}1_{\pi^{\prime}(i)<\pi^{\prime}(j)}\right$. (2)
$\left.\quad+1_{\pi(i)<\pi(j)}1_{\pi^{\prime}(i)>\pi^{\prime}(j)}\right]$
Intuitively, the Kendall-$\tau$ distance counts the differences of all pair-wise comparisons between $\pi$ and $\pi^{\prime}$. For example, let $\pi\;=\;(1,2,3,4)$ and $\pi^{\prime}\;=\;(2,1,4,3)$. Two pairs are discordant among the six unordered pairs: $(1,2),(3,4)$, hence $d(\pi,\pi^{\prime})=2$ and $K_{Mal}(\pi,\pi^{\prime})=\exp(-2l)$.
3 Merge Kernel: Generating Kernels from sorting algorithms
Algorithm 1 Merge feature mapping $\Phi_{Mer}(\pi)$
Input: Permutation $\pi$ of length $n$
Output: Feature vector $\Phi_{Mer}(\pi)$
if length($\pi$) == 1 then
  return []
end if
if length($\pi$) == 2 then
  return [1] if $\pi[0]>\pi$[1] else [0]
end if
 Let $mid=\lfloor\frac{n}{2}\rfloor$
 Let $V_{Left},V_{Right}=\Phi_{Mer}(\pi[:mid]),\Phi_{Mer}(\pi[mid:])$
 Let $\hat{\pi}_{l},\hat{\pi}_{r}=\text{sorted}(\pi[:mid]),\text{sorted}(\pi[mid:])$
 Let $V_{Merge}=[]$, $i=j=0$
while $i<\text{length}(\hat{\pi}_{l})$ and $j<\text{length}(\hat{\pi}_{r})$ do
  if $\hat{\pi}_{l}[i]>\hat{\pi}_{r}[j]$ then
   $V_{Merge}$.append(1)
   $j+=1$
  end if
  if $\hat{\pi}_{l}[i]<\hat{\pi}_{r}[j]$ then
   $V_{Merge}$.append(0)
   $i+=1$
  end if
end while
return $V_{Left}+V_{Right}+V_{Merge}$
In the previous section, we have shown that the core of the Mallows kernel is the pairwise comparison of all elements. Equivalently, it maps a permutation $\pi$ to a feature vector
$\Phi_{Mal}(\pi)\;\in\;\{0,1\}^{\binom{n}{2}}$
where each coordinate corresponds to the comparison of a pair of elements: 0 if they are in ascending order, and 1 otherwise. We then have
$\begin{split}K_{Mal}(\pi,\pi^{\prime})&=\exp\!\Bigl(-\frac{\|\Phi_{Mal}(\pi)-\Phi_{Mal}(\pi^{\prime})\|^{2}}{2\ell^{2}}\Bigr)\\ &=\;K_{\mathrm{RBF}}\bigl(\Phi_{Mal}(\pi),\,\Phi_{Mal}(\pi^{\prime})\bigr).\end{split}$ (3)
under the reparameterization $l=\frac{1}{2\ell^{2}}$ for the length-scale parameter $\ell$ in the Gaussian RBF kernel. Since the RBF kernel $K_{\mathrm{RBF}}$ is strictly positive definite on $\mathbb{R}^{d}$ and thus satisfies Mercer’s condition Mercer (1909), and because positive definiteness is preserved under composition with any deterministic mapping $\Phi$, it follows that $K(\pi,\pi^{\prime})$ constructed from $\Phi$ also satisfies Mercer’s condition and is therefore a valid kernel function.
Thus, other pairwise comparison methods can also be used to construct analogous feature vectors, which—when combined with an RBF kernel—yield valid kernel functions. Naturally, we can extend the idea of pairwise comparison to sorting algorithms: the essence of a sorting algorithm is to compare elements in a sequence and swap them when necessary. Consequently, each sorting algorithm embodies a unique pairwise comparison strategy, suggesting that we can build permutation‐space kernels based on sorting procedures. As a sorting algorithm traverses all elements, it records whether each comparison leads to a swap, thereby fully reconstructing the original permutation; hence, the resulting feature vector retains all information without any loss. Viewed in this light, the Mallows kernel’s exhaustive enumeration of every element pair can be interpreted as an enumeration-sort–inspired featurization, where enumeration sort ranks each item by comparing it with every other element and then places it directly in its final position.
However, not every sorting algorithm can induce a valid feature mapping suitable for kernel construction. This is because the mapping from permutations to feature space must yield feature vectors of fixed length; otherwise, feature vectors of differing lengths would not be compatible with the RBF kernel. Hence, only sorting algorithms that have a fixed comparison path and a constant number of comparisons across all inputs can generate valid feature mappings, e.g., a fixed sorting network with a predetermined comparator sequence Batcher (1968). In fact, most sorting algorithms commonly known to have complexity of order $O(n\log n)$—such as quicksort, heapsort, and standard merge sort—do not satisfy the requirement of constant comparison counts. Nevertheless, merge sort could be an exception: although typically merge sort ceases comparisons once all elements from one subsequence have been merged, redundant comparisons can be artificially introduced during the merge procedure, resulting in a fixed number of comparisons L+R1L+R-1, where $L$ and $R$ are the length of the two subsequences. This ensures that both the comparison path and the number of comparisons remain identical across different permutations, thus establishing a fixed comparison path. Consequently, we select merge sort as the feature mapping $\Phi_{Mer}(\pi)$ for constructing permutation-space kernel function $K_{Mer}$, following Equation 3:
$\begin{split}K_{Mer}(\pi,\pi^{\prime})&=\exp\!\Bigl(-\frac{\|\Phi_{Mer}(\pi)-\Phi_{Mer}(\pi^{\prime})\|^{2}}{2\ell^{2}}\Bigr)\\ &=\;K_{\mathrm{RBF}}\bigl(\Phi_{Mer}(\pi),\,\Phi_{Mer}(\pi^{\prime})\bigr).\end{split}$ (4)
The element-pair comparison mapping $\Phi_{Mer}(\pi)$ is shown in Algorithm 1. Here we present an example for permutation [1,4,3,2] to show the Merging process.
Example: Feature Mapping for $\pi=(1,4,3,2)$
  1. 1.
    Initial Split: $\pi$ splits into $L_{0}=(1,4)$ and $R_{0}=(3,2)$.
  2. 2.
    Recurse on $L_{0}=(1,4)$: Merging sorted lists ‘[1]’ and ‘[4]’ yields feature vector $V_{L}=[0]$.
  3. 3.
    Recurse on $R_{0}=(3,2)$: Merging sorted lists ‘[2]’ and ‘[3]’ yields feature vector $V_{R}=[1]$.
  4. 4.
    Final Merge: Merge sorted lists ‘(1,4)’ and ‘(2,3)’. The fixed comparison path generates the merge vector $V_{Merge}=[0(1<2),1(4>2),1(4>3),1(\text{left padding})]$.
  5. 5.
    Concatenate: The final feature vector is $\Phi_{Mer}(\pi)=V_{L}\oplus V_{R}\oplus V_{Merge}=[0]\oplus[1]\oplus[0,1,1,1]=[0,1,0,1,1,1]$.
We have established that merge sort, with a specially designed fixed-comparison procedure, is uniquely capable of constructing valid kernel functions among common sorting algorithms with a complexity of $\Omega(n\log n)$. We now demonstrate that the feature vector derived from merge sort achieves the theoretical lower bound on vector length for lossless permutation encoding. First, note that the lower bound on time complexity for any comparison-based sorting algorithm is $\Omega(n\log n)$; as this complexity directly corresponds to the number of element comparisons during sorting, it similarly sets a lower bound on the length of the feature vector. On the other hand, consider the permutation space consisting of all $n!$ possible permutations of length $n$. From an information-theoretic viewpoint, encoding all $n!$ permutations without loss using a binary feature vector composed solely of 0,1{0,1} requires a minimum vector length of $\log_{2}(n!)$. Applying Stirling’s approximation Donald et al. (1999), we have: $\log_{2}(n!)=n\log_{2}n-nlog_{2}e+O(\log n)=\Omega(n\log n)$. Consequently, the feature vector length of the merge-sort-based kernel (Merge Kernel) reaches this theoretical lower bound for lossless permutation encoding.
[…]
It is worth noting that, when constructing the Merge kernel via merge sort, we have not required the feature mapping $\Phi$ to possess any group invariance property, such as right-invariance. Traditionally, a permutation-distance measure should be invariant under right multiplication, meaning that applying an identical right-multiplication operation to two permutations should not alter the distance between them. However, only sorting algorithms with complexity $O(n^{2})$ can yield fully right-invariant kernel functions with no information loss (otherwise, a simple Spearman’s footrule Diaconis & Graham (1977) with $O(n)$ complexity can hold right-invariance as well), since such invariance necessitates exhaustive pairwise comparisons among all $\frac{n(n-1)}{2}$ pairs of elements—an impossibility for more efficient sorting algorithms like merge sort. Consequently, although the Merge kernel achieves better computational efficiency through a more compact encoding, it sacrifices a certain degree of performance due to the loss of right-invariance.
4 Experiments
4.1 Benchmarks and Experiment Settings
Our empirical evaluation focuses on the state-of-the-art BOPS-H algorithm Deshwal et al. (2022) as the sole baseline. This choice is twofold: first, BOPS-H was shown to substantially outperform other permutation-specific methods like COMBO Oh et al. (2019). Second, our core objective is a principled comparison between our Merge Kernel and the Mallows Kernel—both natively designed for permutations. We therefore exclude methods for high-dimensional continuous BO (e.g., TuRBO, REMBO) to ensure a direct evaluation of representation power without confounding factors from domain adaptation. To comprehensively test our scalability hypothesis, we evaluate both methods on a suite of low- and high-dimensional benchmarks.
4.1.1 Low-dimensional Benchmarks
The study by Deshwal et al. (2022) exclusively considers problems with dimensions of 30 or less. Accordingly, we adopt their experimental settings to form our suite of low-dimensional benchmarks. The BOPS-H algorithm follows and modifies the local-search strategy used in COMBO, examining only the set of neighbouring permutations of the current incumbent to restrict the combinatorial search space, we therefore adopt the same procedure in our experiments. GPyTorch Gardner et al. (2018) and BoTorch Balandat et al. (2020) libraries are used to implement both algorithms. Expected Improvement acquisition function is used for all the experiments, and 10 restarts are used for local search based EI optimization for BOPS-H and MergeBO. Each benchmark is evaluated with 20 independent trials, each consisting of 50 iterations. The random seed for each trial is set to its trial index.
We evaluate our method on the same two synthetic benchmarks and the same two real-world applications in Deshwal et al. (2022). Detailed information for all benchmarks are listed below:
(1)Quadratic Assignment (QAP)n=15. This is a classic facility-location problem: assign $n$ facilities to $n$ locations so that flow costs and distances align optimally. We use the 15-city instances from QAPLIB Burkard et al. (1997), each with cost matrix $A$ and distance matrix $B$, and minimise $\operatorname{Tr}(APBP^{\mathsf{T}})$ over permutation matrices $P$.
(2)Travelling Salesman (TSP)n=15. The TSP seeks the shortest Hamiltonian cycle through a set of cities and is a standard benchmark for route-planning. Our instances are 15 city PCB drill tours from TSPLIB Reinelt (1995); the score is the total travel time to visit all holes exactly once and return.
(3)Floor Planning (FP)n=30. Floor planning is an NP-hard VLSI layout task that packs rectangular modules on a chip while minimising area and manufacturing cost. We evaluate two 15-block variants (FP-1 and FP-2); each permutation defines a block placement whose cost we minimise.
(4)Cell Placement (CP)n=30. Cell placement arranges logic cells on a row to reduce wire-length and hence circuit delay. We consider 30 equal-height cells with a fixed net-list; the objective is the total Manhattan wire-length induced by a permutation of cell positions.
Because the publicly available implementation of the Mallows kernel (https://github.com/aryandeshwal/BOPS) does not provide the interface required to run the Rodinia’s heterogeneous many-core benchmark Che et al. (2009), we did not perform experiments on this benchmark. We note a numerical discrepancy between our replicated results and those reported in the original paper, which we attribute to subtle implementation details not specified in the publication, such as problem instance choices. However, we emphasize that within our experimental framework, both the Merge Kernel and the Mallows Kernel were evaluated under identical conditions, ensuring a fair and controlled comparison of their relative performance.
4.1.2 High-dimensional Benchmarks: Traveling Thief Problems
We introduce traveling thief problems (TTP) Bonyadi et al. (2013) as high-dimensional benchmarks, which is a defined as a combination of TSP and knapsack problem: a thief must determine a tour through $n$ cities with distance matrix $D=\{d_{ij}\}$ while simultaneously selecting items of varying weights $w_{k}$ and values $p_{k}$ to maximize profit without exceeding a knapsack capacity $W$. This structure defines a complex, hybrid search space, combining an $n$-dimensional permutation space for the city tour with a $\{0,1\}^{m}$ discrete space for item selection. Despite its typical application in evaluating white-box or heuristic algorithms Polyakovskiy et al. (2014); Gupta et al. (2015), we adapt the TTP as a true black-box benchmark, providing no structural information to the optimizer.
Following the implementation in Polyakovskiy et al. (2014), we create three distinct instances based on a $n=280$-city problem (a 280-dimensional permutation space). These instances feature a large number of items with varying properties: (1) $m=279$ items with uncorrelated weights; (2) $m=837$ items with bounded strong correlation in the weights; and (3) $m=837$ items with uncorrelated weights. These benchmarks provide a strenuous test for our proposed MergeBO and the baseline BOPS-H. Notice that the TTP is a hybrid space problem, we modified both MergeBO and BOPS-H by multiplying an RBF kernel on the $\{0,1\}^{m}$ discrete space:
$K_{TTP}((\pi,\sigma),(\pi^{\prime},\sigma^{\prime}))=K(\pi,\pi^{\prime})K_{RBF}(\sigma,\sigma^{\prime})$ (5)
Where $\sigma$ is the item picking strategy. The neighbouring permutation search method for BOPS-H is computationally infeasible on such a vast space. Therefore, we adopt a continuous relaxation approach, treating the entire set of optimization variables $(\pi,\sigma)$ as a continuous vector for gradient-based optimization, and the result is subsequently projected back to the nearest permutation and binary vectors.
All experiments use the EI acquisition function. Each benchmark is evaluated with 50 independent trials, each consisting of 55 iterations (5 iterations of random initialization and 50 iterations of optimization). The random seed for each trial is set to its trial index.
4.1.3 Evaluation Metrics
In this study, we employ two evaluation metrics: the final simple regret and the area under the best‑so‑far regret curve (AUC). In optimization, regret is defined as the difference between the best objective value observed to date and the global optimum:
$r_{t}=f_{t}^{\text{best}}-f^{*}$
Following this concept, the final simple regret is the regret value obtained in the last iteration and reflects the algorithm’s ultimate optimization capability when computational cost is disregarded. By contrast, the regret AUC, or cumulative regret—the sum (area under the curve) of the simple regret across all iterations—quantifies the convergence speed of the entire optimization process:
$\text{AUC}_{T}=\Sigma r_{t}$
Generally speaking, the final simple regret and AUC represent different, valuable aspects of an algorithm’s performance: the former represents the the quality of the solution it ultimately finds while the latter represents the “journey”, or convergence speed during optimization. For sample-efficient methods like BO, these two evaluation metrics are standard practices as they allow for a comprehensive comparison.
Problem Dimension Merge feature length Mallows feature length
TSP 15 45 105
QAP 15 45 105
FP 30 119 435
CR 30 119 435
TTP 280 2009 39060
Table 1: Feature length comparison of Merge and Mallows kernel over problems of different scales.
It is important to note that this study does not include a comparison of wall-clock computation times. This is a deliberate choice grounded in the fundamental premise of BO, where the cost of function evaluations (e.g., physical experiments or complex simulations) is assumed to far outweigh the computational cost of the algorithm itself. Consequently, our analysis prioritizes metrics related to sample efficiency, which is the primary bottleneck in such real-world scenarios. As a more stable and implementation-agnostic proxy for computational complexity, we instead report the feature vector dimensions generated by each kernel in Table 1, which directly reflects the compactness of the learned representations. Furthermore, as our experiments were conducted on a shared high-performance computing (HPC) cluster, reported wall-clock times would be subject to scheduler-induced variability, making them an unreliable metric for rigorous algorithmic comparison.
4.2 Results and Discussion
Figure 1: Low- and high-dimensional results comparing Mallows kernel (BOPS-H) and Merge kernel (MergeBO) on the current regret value (difference between best-so-far and optimal value) vs. number of iteration. Solid lines show the average regrets, while the shaded areas denote one standard deviation.
Simple final regret Best so far AUC
Merge Mallows Merge Mallows
[…]
Simple final regret Best so far AUC
Merge wins Ties Mallows wins Merge wins Ties Mallows wins
$\text{TSP}_{n=15}$ 1 12 7 5 0 15
$\text{QAP}_{n=15}$ 1 3 16 1 2 17
$\text{FP}_{n=30}$ 10 4 6 10 1 9
$\text{CR}_{n=30}$ 9 3 8 8 0 12
$\text{TTP1}_{n=280}$ 50 0 0 50 0 0
$\text{TTP2}_{n=280}$ 50 0 0 50 0 0
$\text{TTP3}_{n=280}$ 50 0 0 50 0 0
Table 2: Performance comparison between MergeBO and BOPS-H (Mallows kernel), bold font indicates statistically significant superiority as determined by a binomial sign test (p << 0.05, corresponding to more than 15 wins of 20 trials).
Our experimental results, as presented in Table 2 and Figure 1, systematically reveal a strong correlation between the performance advantage of the Merge kernel and the problem’s dimensionality. For low-dimensional permutation problems (TSP and QAP), the Mallows kernel, which is specifically designed for such tasks, exhibited a slight performance advantage. However, on problems of intermediate dimensionality (FP and CR), the performance of the two kernels was comparable, with no statistically significant difference observed. This trend shifted decisively on the higher-dimensional TTP problems, where the Merge kernel demonstrated definitive superiority. The statistical data in Table 2 indicates that the Merge kernel significantly outperformed the Mallows kernel across all TTP instances on both the final regret and AUC metrics. Furthermore, the convergence curves in Figure 1 confirm its substantially faster convergence speed. This validates the superior performance and scalability of our proposed method on complex, high-dimensional optimization problems.
These results corroborate our core hypothesis: the Merge kernel possesses an inherent advantage in high-dimensional permutation optimization problems, owing to its more compact structural design. The disparity between its $O(n\log n)$ feature complexity and the Mallows kernel’s $O(n^{2})$ complexity widens dramatically as the dimensionality n increases. This is explicitly quantified by the feature length comparison in Table 1: for low, intermediate, and high-dimensional problems, the feature dimensionality of the Mallows kernel is approximately 2, 3.5, and 19.5 times that of the Merge kernel, respectively. These results indicate that the experimental performance is governed by a trade-off between two key factors: (1) the ability to capture global information via the distance metric, and (2) the search efficiency driven by the compactness of the feature space. Due to its right-invariance property, the Mallows kernel possesses a stronger distance metric capability than the Merge kernel. However, as dimensionality increases, the vast disparity in feature vector length leads to a more pronounced space-compression effect. The resulting gains in search efficiency begin to outweigh the performance benefits afforded by the superior distance metric. Consequently, as the problem dimensionality continues to grow, the performance of the Merge kernel ultimately surpasses that of the Mallows kernel by a significant margin. This naturally suggests a dimension-dependent heuristic for practitioners: leveraging the Mallows kernel’s robust, right-invariant distance metric for low-dimensional tasks, while switching to the Merge kernel to capitalize on its superior scalability in high-dimensional regimes.
5 Conclusions
In this work, we proposed a novel kernel construction framework for permutation spaces by leveraging sorting algorithms as structured comparison schemes. Within this framework, we introduced the Merge kernel—an efficient, compact, and theoretically grounded alternative to the quadratic Mallows kernel. We showed that the Merge kernel achieves the information-theoretic lower bound on feature complexity while preserving meaningful structural information. This contribution bridges sorting theory and kernel design, revealing a fundamental trade-off between a model’s structural invariance and the compactness of its feature space.
Empirical experiments confirmed this trade-off: while the state-of-the-art BOPS-H algorithm held a marginal advantage in low-dimensional problems, their performances were comparable on intermediate-dimensional tasks. In high-dimensional settings, however, MergeBO’s compact representation enabled far superior search efficiency, allowing it to significantly outperform BOPS-H. This work opens exciting possibilities for scaling Bayesian optimization to larger permutation spaces. Future directions include exploring other sorting algorithms with constant comparison counts, and applying this framework to challenging real-world domains like combinatorial neural architecture search and computational biology.
Appendix A The Use of Large Language Models
We declare that the large language models (LLMs) ChatGPT and Gemini are only used to aid and polish writing. No further applications of LLMs are used in this research, including but not limited to retrieval, research ideation and experiment designing.